OpenStudy (anonymous):

what is the area and width of a triangle with the measures of 29.6=21+2w

OpenStudy (amistre64):

i take it that that is the perimeter of the triangle in question...

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

make s = the area is half the perimeter then use Herons formula sqrt(s(s-a)(s-b)(s-c)) = area. if I remember it correctly

OpenStudy (amistre64):

that 2w seems a bit off, I would assume that means that 2 of the legs are of equal length and that they measure "w" in length

OpenStudy (amistre64):

maybe it means after you find the area, use what you know for "w" to plug into the normal area formula. Area = width(height)/2

OpenStudy (anonymous):

can you answer this one..... what is the area and width of a rectangle with the perimeter of 29.6 and the length of 10.5..? i used P=2l+2w but i dont know how to solve it

OpenStudy (amistre64):

29.6 - 21 = 8.6 which dont make no sense, side thats 21, you would need 2 more sides that add up to at least 21. All thats left here is 17.2. Im thinking that either theres imformation missing or that this dont make a triangle

OpenStudy (amistre64):

but then again, im using my georgia math....

OpenStudy (anonymous):

well can you answer this... Use P=(2)l+(2)w so far I have 20.6=(2)l+(2)3.1 can you solve that so i know the width?